Please written down the answer and formula on paper
Stat 50 Final Exam Last Name- First Name - The population standard deviation for the age of Valley College students is 15 years. If we wan population mean age of Valley College students, how many randomly selected Valley College students must be surveyed? 7. t to be 99% confident that the sample mean age is within two years of the true A common claim is that garlic lowers cholesterol levels. In a test of the effectiveness of garlic, 97 subjects were treated with doses of raw garlic, and their cholesterol levels were measured before and after the treatment. The changes in their levels of LDL cholesterol (in mg/dL) have a mean of 0.5 and a standard deviation of 17.0. 8. ー Use the sample statistics of nーーーーー construct a 98% confidence interval estimate of the mean net change in LDL cholesterol after the garlic treatment a. ーーーーーand s-- to b. What does the confidence interval suggest about the effectiveness of garlic in reducing LDL cholesterol?
Answer #1

7)

Confidence level = 0.99

Zc = 2.58 ( Using z table)

Margin of error = E = 2

Population standard deviation = sigma = 15

We have to find sample size (n)

n=left ( rac{Zc*sigma}{E} ight )^{2}

2.58 15 -(2.38.15) rl

n = 19.352-374

--------------------------------------------------------------------------------------------------------------------------------------

8)

Sample size = n = 97

Sample mean = ar{x} = 0.5

Standard deviation = s = 17

a) We have to construct 98% confidence interval.

Formula is

(ar{x}-E , ar{x}+E)

Here E is a margin of error.

E=

Degrees of freedom = n - 1 = 97 - 1 = 96

Level of significance = 0.02

tc = 2.366

2.366 17 4.0836 97

So confidence interval is ( 0.5 - 4.0836 , 0.5 + 4.0836) = > ( -3.5836 , 4.5836)

b) Since the confidence interval is wider this implies that there are larger chance of obtaining an observation within that interval. Hence, we could say that the use of garlic lowers LDL cholesterol.